Could someone please double check my work please,
f. What is the profit made from selling 20 tiles per month?(answer- worth 3 points - must show work)
P(20)=-20^2+(33*20)-200
P(20)=400+660-200
P(20)=460
The profit made from selling 20 tile sets per month is $460.00
g. What is the profit made from selling 25 tiles each month?(answer- worth 3 points - must show work)
P(25)=-25^2+(33*25)-200
P(25)=625+825-200
The profit made from selling 25 tile sets is $1250.

Answer by ewatrrr(24785)

 
Hi,
Re: TY Note: Graph demonstrates Profit Peaks at ~16 tiles and then decreases as more tiles are sold
Careful, the (-1) stays in effect
P(x)= (-1)x^2 + 33x - 20
P(20)= (-1)20^2+(33*20)-200
P(20)= -400+660-200 = 60
P(25)= (-1)25^2+(33*25)-200
P(25)= - 625+825-200 = 0